I need to figure out the third side of a right triangle. I do not know how to figure it out.
The bottom of the right triangle is "a". The left = 1 and the slant of the triangle = sqrt 5

Answer by stanbon(75887)

I need to figure out the third side of a right triangle. I do not know how to figure it out.
The bottom of the right triangle is "a". The left = 1 and the slant of the triangle = sqrt 5
--------------------------
hypotenuse^2 = leg^2 + leg^2
(sqrt(5))^2 = a^2 + 1^2
5 = a^2 + 1
a^2 = 4
a = 2
-------------
Cheers,
Stan H.